OpenStudy (anonymous):

Solve the following system of equations. Enter the y-coordinate of the solution. Round your answer to the nearest tenth. 5x+2y=7 -2x+6y=9

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

Thanks.

OpenStudy (radar):

Try x=12/17 y=177/102 Do the rounding.